OpenStudy (anonymous):

Find the distance between the points (3, 8) and (-1, 9)

OpenStudy (anonymous):

\[\sqrt{5}\] \[\sqrt{15}\] \[\sqrt{17}\]

OpenStudy (anonymous):

Answers^

OpenStudy (anonymous):